Determine whether u=(2,8,2) is a linear combination of u1=(2,-2,0) u2=(3,0,-3) and u3=(-2,0,-1) or not


1
Expert's answer
2021-08-09T16:19:30-0400

"solution\\
given \space 
\\
u=(2,8,2) \space \\
 \space u_1=(2,-2,0), \space u_2=(3,0,-3) \space and \space u_3=(-2,0,-1) \space \\
now \space we \space will \space be \space check \space u \space is \space a \space linear \space combination \space of \space u_1, \space u_2 \space and \space u_3 \space or \space not,\\
first \space of \space all \space we \space write \space a \space linear \space combination \space form \space of \space u \space in \space terms \space of \space \space u_1, \space u_2 \space and \space u_3\\
such \space that\\
u=a.u_1+b. \space u_2 \space +c.u_3\\
(2,8,2) \space =a.(2,-2,0)+b. \space (3,0,-3) \space +c.(-2,0,-1)...(A) \space \\
now \space we \space solve \space this \space expression \space to \space find \space a \space ,b \space and \space c \space \\
if \space we \space will \space be \space get \space value \space of \space a \space ,b \space and \space c \space \space then \space u \space is \space a \space linear \space combination \space of \space u_1, \space u_2 \space and \space u_3 \space otherwise \space not\\
---------------------------------------\\
again\\
(2,8,2) \space =a.(2,-2,0)+b. \space (3,0,-3) \space +c.(-2,0,-1) \space \\
(2,8,2) \space =(2a+3b-2c,-2a,-3b-c)\\
hence\\
2a+3b-2c=2...(1)\\
-2a=8...(2)\\
-3b-c=2...(3)\\
from \space (2)\\
a=-4...(4)\\
put \space value \space of \space a \space in \space (1)\\
3b-2c=10...(5)\\
now \space solve \space (5) \space and \space (3), \space we \space get \space \\
b=\frac{2}{3}, \space and \space c=-4...(6)
---------------------------------------\\
 \space we \space get \space \\
a=-4, \space b=\frac{2}{3}, \space and \space c=-4\\
hence \space we \space say \space \\
 \space u \space is \space a \space linear \space combination \space of \space u_1, \space u_2 \space and \space u_3 \space \\
put \space value \space of \space a \space ,b \space and \space c \space \space in \space eq(A)\\
we \space get\\
(2,8,2) \space =(-4).(2,-2,0)+(\frac{2}{3}). \space (3,0,-3) \space +(-4).(-2,0,-1)\\"